OpenStudy (anonymous):

Write an equation in slope-intercept form for each line through (3, 12) and (6, 27)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(3\quad ,&12)\quad &(6\quad ,&27) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"} \)